a progressive and stationary simple hamonic wave having frequency 250hz and each having same velocity 30m/s 1) determine phase difference between two vibrating points 2) wave equation of progressive wave if amplitude is 0.03m 3) distance between nodes in stationary wave

Grade:12th pass

1 Answers

Arun
4 years ago
period = T = 1/f = 1/250
wavelength = L = speed*time = 30/250

sin 2pi(t/T-x/L)
x/L = .10 (250/30)
phase = 2 pi (.1*250/30)

half a wavelength between nodes = 15/250